An airplane flying with the wind can cover 600 miles in two hours. The return trip against the wind takes two and one half hours. What is the speed of the plane in still air?
Answer by mananth(16949)

let speed of plane in still air be x
let speed of wind be y
..
speed with wind = x+y
2(x+y)=600
..
speed against wind = x-y
2.5(x-y)=600
multiply by 10
25x-25y= 6000( multiply by 2)
2x+2y=600 ( multiply by 25)
ADD
..
50x-50y+50x+50y= 12000 +15000
100x=27000
x=270 mph speed in still air